OpenStudy (malcolmmcswain):

Three consecutive integers add up to 51. Which integers are they?

OpenStudy (tkhunny):

Define three consecutive integers.

OpenStudy (malcolmmcswain):

Integers that follow each other. I.E. 1,2,3

OpenStudy (texaschic101):

3 consecutive integers.. x, x + 1, x + 2 they total 51... x + (x + 1) +( x + 2) = 51 now, combine like terms and solve for x

OpenStudy (texaschic101):

and once you find x, then sub that back into x + 1 = and x + 2 = and that will give you all 3 numbers

OpenStudy (malcolmmcswain):

3x+3=51

OpenStudy (texaschic101):

now solve for x

OpenStudy (malcolmmcswain):

3x=51-3 3x=48

OpenStudy (malcolmmcswain):

x=16

OpenStudy (malcolmmcswain):

So, 16, 17, 18.

OpenStudy (texaschic101):

check...add them and see if they equal 51

OpenStudy (malcolmmcswain):

16+17+18 = 51

OpenStudy (malcolmmcswain):

Thanks!

OpenStudy (texaschic101):

sure thing :)